OpenStudy (anonymous):

Your friend drops your house keys down to you from their second floor window. If you catch them 2.4 m from where your friend dropped them, what is the velocity of the keys when you catch them?

OpenStudy (anonymous):

a. -6.9 m/s c. -7.2 m/s b. -6.5 m/s d. -7.4 m/s

OpenStudy (aaronq):

the keys are accelerating downwards at 9.8 m/s^2 but have no initial velocity (they were dropped not thrown). Always draw a diagram  \(y_i=2.4~ m\) \(y_f=0~ m\) \(v_i=0\) use: \(v_f^2=v_i^2+2a\Delta x\) \(v_f^2=(0)^2+2(-9.8~m/s^2)(0~m-2.4~ m)\) \(v_f=\sqrt{47.04}=6.8585~m/s \approx 6.9~m/s\)